Problem 6-33 (LO. 5, 6, 8) Five years ago, Gerald invested $150,000 in a passive activity, his sole investment venture. On January 1, 2018, his amount at risk in the activity was $30,000. His shares of the income and losses were as follows:

Year Income (Loss)

2018 ($40,000)

2019 (30,000)

2020 50,000

Gerald holds no suspended at-risk or passive activity losses at the beginning of 2018. If an answer is zero, enter "0".

a. If losses were limited only by the at-risk rules, how much can Gerald deduct in 2018 and 2019? Year Loss Allowed Suspended 2018 ($40,000) $ $ 2019 ($30,000) Total $

b. Refer to the information in part (a) above. If losses were limited by the at-risk and the passive activity loss rules, how much would Gerald be able to deduct in 2018 and 2019? Of the allowable at-risk loss for 2018, Gerald may deduct $ in 2018 due to the passive loss rules and of the allowable at-risk loss from 2019, Gerald may deduct $ in 2019 due to the passive loss rules.

c. Assuming Gerald has $50,000 income in 2020, (and considering both at-risk and passive activity loss rules), what is the amount of Gerald's suspended losses at the end of 2020? Suspended under the at-risk rules: $

Suspended under the passive activity loss rules: $

What is his taxable income for 2020? $

At the end of 2020, what is the amount of Gerald's adjusted basis in the activity? $
Answer #1

Part A

Year

Loss

Allowed

Suspended

2018

(40000)

30000

10000

2019

(30000)

0

30000

Total carryover to 2020

40000

Due to amount at risk in the activity, the deduction is limited to $30000 in 2018

No deduction in 2019 as there is no amount at risk in the activity

Part B

2018

$0

2019

$0

Of the allowable at-risk loss for 2018, Gerald may deduct $0 in 2018 due to the passive loss rules and of the allowable at-risk loss from 2019, Gerald may deduct $0 in 2019 due to the passive loss rules.

Part C

Suspended under the at-risk rules

$0

Suspended under the passive activity loss rules

$0

Balance of suspended losses at 2020 = $50,000 income increases Gerald's at-risk amount to $50,000 - disallowed at-risk losses from 2018 and 2019 ($30,000 + $10,000) that can be used in 2020 = 50000-40000 = 0

Part D

Taxable income

$30000

Amount of Gerald's adjusted basis in the activity

$10000

Amount of Gerald's adjusted basis in the activity = 50,000 income increases Gerald's at-risk amount to $50,000 - disallowed at-risk losses from 2018 and 2019 ($30,000 + $10,000) that can be used in 2020 = 50000-40000 = $10000